Person is jogging at 7m/s and then accelerated for 2 seconds covering a distance of 60 meters. What is the acceleration?
1
Expert's answer
2020-09-14T10:20:21-0400

The displacement for accelerated motion is "S(t) = v_0 t + \frac{a t^2}{2}", where "v_0" is initial speed and "a" is acceleration. From the last formula, the acceleration is "a = \frac{2(S - v_0 t)}{t^2} = \frac{2 (60 m - 7\frac{m}{s} \cdot 2 s)}{4 s^2} = 23 \frac{m}{s^2}".